### v1

# Bound excited states of Fröhlich polarons in one dimension

Abstract
The one-dimensional Fröhlich model describing the motion of a single electron interacting with optical phonons is a paradigmatic model of quantum many-body physics. We predict the existence of an arbitrarily large number of bound excited states in the strong coupling limit and calculate their excitation energies. Numerical simulations of a discretized model demonstrate the complete amelioration of the projector Monte Carlo sign problem by walker annihilation in an infinite Hilbert space. They reveal the threshold for the occurrence of the first bound excited states at a value of $\alpha\approx 1.73$ for the dimensionless coupling constant. This puts the threshold into the regime of intermediate interaction strength. We calculate the spectral weight and the number density of the dressing cloud in the discretized model.
Studies of the polaron, describing the motion of an electron in a polarizable medium, go back almost a century to the works of Landau [1] and Pekar [2]. Formalised by Fröhlich [3], the polaron model is one of the most studied problems of quantum many-body theory [4] with real-world relevance for electronic material properties [5]. Calculations of the ground state energy, the dispersion relation and also the structure of the excitation continuum have been achieved at great accuracy and become an important test bed for analytical [6] and numerical [7, 8, 9, 10] techniques. Thus it may seem surprising that evidence for bound excited states below the excitation continuum has emerged only very recently [11].
The proof of the existence of arbitrarily many bound excited states in the strong-coupling limit of the three-dimensional model [11] demonstrates that previous approaches have missed key features of the Fröhlich polaron at intermediate and strong coupling. Previous numerical studies have failed to produce evidence for bound excited states in the three-dimensional case [8]. Studies of one-dimensional Fröhlich polaron models found excitations of the electron as resonances at high energies within the phonon continuum [12, 13]. In contrast, for the Holstein model of a small polaron, the existence of a bound excited state below the continuum is well known [14, 15, 16].
In this Letter we calculate the energies of the bound excited states of the phonon dressing cloud for the Fröhlich model in one dimension in the strong coupling limit, and provide numerical evidence for them in the weak to intermediate coupling regime applying Full Configuration Interaction Quantum Monte Carlo (FCIQMC) [17, 18, 19]. The first bound excited state appears at $\alpha\approx 1.73$, which is in a crossover regime between weak and intermediate coupling strength with a spectral weight of $\approx 0.2$, suggesting that these states may be accessible in experiments.
FCIQMC is a modern flavor of Projector Quantum Monte Carlo [20]. Originally developed for electronic structure calculations of molecules [17], it has been extended to bosonic systems [21, 22] and electron-phonon systems with infinite Hilbert space [23]. Sampling the ground and excited [19, 24] states of the Fröhlich model in Fock space provides us with direct access to the energies and properties of these states, avoiding the ill-conditioned analytic continuation procedure required to obtain spectral information from diagrammatic Monte Carlo calculations [8]. Our state-of-the art semistochastic [18, 25] implementation of FCIQMC [26, 27] allows us to sample ground and excited states with moderate stochastic error and well-controlled systematic biases. Specifically, we report the first successful amelioration of the sign problem in Projector Monte Carlo [28] by walker annihilation in an infinite Hilbert space without employing sign-problem-mitigating approximations. Previously, this was only achieved in calculations with a finite Hilbert space [17, 29, 30, 31].
The model – The one-dimensional Fröhlich model describes an electron interacting with lattice vibrations in one dimension. The original three-dimensional Fröhlich Hamiltonian [3] was generalized to arbitrary dimensions in Refs. [32, 33]. The Hamiltonian in the one-dimensional case is
$\hat{H}$ $=\frac{\hat{p}^{2}}{2m}+\hbar\omega_{\mathrm{LO}}\sum_{k}\hat{a}^{\dagger}_{k}\hat{a}_{k}$
$-\hbar\omega_{\mathrm{LO}}\sqrt{\frac{2\alpha}{L/l_{0}}}\sum_{k}\left(\hat{a}_{k}e^{ik\hat{r}}+\hat{a}^{\dagger}_{k}e^{-ik\hat{r}}\right),$ (1)
where $\hat{p}$ and $\hat{r}$ are the electron momentum and position, $\hat{a}^{\dagger}_{k}$ and $\hat{a}_{k}$ are the phonon creation and annihilation operators, $\omega_{\mathrm{LO}}$ is the frequency of longitudinal optical phonons (taken to be constant), $m$ is the band mass of the electron, and $\alpha$ is a dimensionless coupling constant. $L$ is the length of the system with periodic boundaries (eventually taken to be infinite), and $l_{0}=\sqrt{{\hbar}/{2m\omega_{\mathrm{LO}}}}$ is a natural unit of length.
The energy-momentum spectrum consists of an isolated ground state band [34] with decreasing energy as the coupling strength increases, and a continuum of scattering states starting at an energy of $1\hbar\omega_{\mathrm{LO}}$ above the ground state [35]. Although the absence of discrete excited states can be proven at weak coupling [36], discrete (bound) excited states with energies below the continuum exist at intermediate and strong coupling, as will be shown in the following.
Excitation energies in the strong coupling limit – For large $\alpha$, the discrete excited eigenvalues of $\hat{H}$ in Eq. (Bound excited states of Fröhlich polarons in one dimension) at total momentum $P=0$ can be obtained analytically, and are given by
$\frac{E_{j}}{\hbar\omega_{\mathrm{LO}}}\approx-\frac{\alpha^{2}}{3}-\frac{1}{2}\sum_{i\geq 0}(1-\omega_{i})+\omega_{j},\quad j=0,1,\dots$ (2)
As is well-known [37, 38, 39], the first term, $-\alpha^{2}/3$, represents the semiclassical energy obtained by minimizing the corresponding Pekar functional with respect to the electron wave function $\psi(x)$ and the classical polarization field $\phi(x)$. The Pekar functional (given in dimensionless units integrating over $x=r/l_{0}$)
$\mathcal{E}^{\rm P}[\psi,\varphi]=\int|\psi^{\prime}|^{2}-2\sqrt{2\alpha}\int|\psi|^{2}\operatorname{Re}(\varphi)+\int|\varphi|^{2}$ (3)
is obtained as an expectation value of the Hamiltonian of Eq. (Bound excited states of Fröhlich polarons in one dimension) with respect to a product of the electron wave function $\psi(x)$ and a coherent state for the phonon field with amplitude $\varphi(x)$. The physical picture is that, at large coupling, the polarization field effectively becomes classical, and the electron occupies the ground state in the corresponding effective potential. The Pekar functional has the explicit minimizer $\psi^{\rm P}(x)=\sqrt{\alpha/2}\,\mathrm{sech}(\alpha x)$ for the electron wave function and $\varphi^{\rm P}(x)=\sqrt{2\alpha}\psi^{\rm P}(x)^{2}$ for the polarization field. The other terms in Eq. (2) involving the ($\alpha$ independent) frequencies $\omega_{j}$ quantify the quantum fluctuations of the field around its classical value [40, 41, 42, 43, 44, 11]. Letting the electron take the instantaneous ground state in a varying field, they are obtained by expanding $\min_{\psi}\mathcal{E}^{\rm P}[\psi,\varphi^{\rm P}+{\varphi}]$ to second order in ${\varphi}(x)$, with the result
$-\frac{\alpha^{2}}{3}+\|\text{Im}({\varphi})\|^{2}+\langle\text{Re}({\varphi}),\hat{\omega}^{2}\text{Re}({\varphi})\rangle.$ (4)
Here the Hessian operator takes the simple form
$\hat{\omega}^{2}=1+8\alpha\partial\psi^{\rm P}(-\partial^{2}+\alpha^{2})^{-2}\psi^{\rm P}\partial,$ (5)
where $\psi^{\rm P}(x)$ is understood as a multiplication operator, and $\partial=d/dx$. By scaling, the eigenvalues of $\hat{\omega}$ are independent of $\alpha$. We denote them by $10=\omega_{0}<\omega_{1}\leq\dots<$ 1. There is a single zero mode, $\omega_{0}=0$, resulting from translation invariance. The lowest non-vanishing eigenvalues are given by $\omega_{1}\approx 0.647$ and $\omega_{2}=\sqrt{7/9}\approx 0.882$. For details, we refer to Appendix A in the End Matter.
We proceed by quantizing the field $\varphi(x)\to\hat{\varphi}(x)$. Expanding $\hat{\varphi}(x)=\sum_{j\geq 0}\hat{c}_{j}u_{j}(x)$ in terms of the (real-valued) eigenfunctions $u_{j}(x)$ of $\hat{\omega}^{2}$, and defining $\hat{p}_{j}=\sqrt{2}\text{Im}(\hat{c}_{j})$ and $\hat{q}_{j}=\sqrt{2}\text{Re}(\hat{c}_{j})$, we re-interpret the quadratic terms in Eq. (4) as a sum of harmonic oscillators with frequencies $\omega_{j}$:
$\frac{1}{2}\sum_{j\geq 0}\left(\hat{p}_{j}^{2}+{\omega_{j}^{2}}\hat{q}_{j}^{2}\right).$ (6)
This leads to a contribution to the ground state energy of the form $\frac{1}{2}\sum_{j\geq 0}\omega_{j}=\infty$ from the zero-point motion. To obtain the finite quantum corrections, we need to subtract the (infinite) energy of the fluctuations in the absence of the electron, however, leading to the corrected expression $\tfrac{1}{2}\sum_{j\geq 0}(\omega_{j}-1)$, hence justifying Eq. (2) for $j=0$. Numerically, one finds $\tfrac{1}{2}\sum_{j\geq 0}(\omega_{j}-1)\approx-0.95$. For $j\geq 1$, the excitation energies correspond to excitations of different oscillators in Eq. (6). Since $\omega_{j}>\tfrac{1}{2}$ for $j\geq 1$, only single excitations can appear below the continuum. The zero mode $j=0$ needs to be discarded for the calculation of the excitation spectrum of Eq. (Bound excited states of Fröhlich polarons in one dimension) due to the restriction to zero total momentum.
Lee-Low-Pines transformation – Using a unitary transformation due to Lee, Low, and Pines amounting to a frame transformation [45], the electron coordinate dependence of the Fröhlich polaron Hamiltonian of Eq. (Bound excited states of Fröhlich polarons in one dimension) can be eliminated and the Hamiltonian rewritten in terms of the conserved total momentum $P$:
$\hat{H}_{\mathrm{LLP}}$ $=\frac{1}{2m}\left(P-\hbar\sum_{k}k\hat{a}^{\dagger}_{k}\hat{a}_{k}\right)^{2}+\hbar\omega_{\mathrm{LO}}\sum_{k}\hat{a}^{\dagger}_{k}\hat{a}_{k}$
$-\hbar\omega_{\mathrm{LO}}\sqrt{\frac{2\alpha}{L/l_{0}}}\sum_{k}\left(\hat{a}^{\dagger}_{k}+\hat{a}_{k}\right).$ (7)
This equivalent Hamiltonian acts only on a Fock space of phonons for a given value of the total momentum $P$ and is more convenient for numerical calculations. We take the system size $L$ finite, which discretizes the phonon modes with the values of $k$ restricted to multiples of $2\pi/L$. A finite number of modes $M$ used in numerical calculations results in a momentum cutoff of $k_{c}=\pi(M-1)/L$. Even so, the phonon Hilbert space dimension is still formally (and practically) infinite, as an arbitrary number of phonons can reside in each mode. Technically, the number of phonons in each mode is limited to be $<256$ by the 8-bit number type used for representing the Fock basis. We perform calculations with up to $M=73$ and Hilbert space dimension $256^{M}\approx 6\times 10^{175}$. In the following, we will restrict the analysis to $P=0$.
FCIQMC – The algorithm samples the ground state of $\hat{H}_{\mathrm{LLP}}$ by repeatedly applying the operator $1-\text{d}\tau(\hat{H}_{\mathrm{LLP}}-S)$ to a coefficient vector. The shift parameter $S$ is updated to keep the norm (the number of ‘walkers’) approximately constant [21], and therefore provides an estimator of the eigenstate energy after equilibration. The operation is performed exactly for coefficients with a modulus above a given threshold and stochastically for the rest. Combined with stochastic vector compression [46], this implements a semistochastic algorithm, where part of the Hamiltonian matrix is treated deterministically while the rest is sampled stochastically [18, 25]. Crucially, contributions of opposite sign to a coefficient can annihilate and thus reduce sign noise. At sufficiently large walker number, the sign of the dominant coefficients becomes stable by a spontaneous symmetry-breaking process [28, 47] and the sign problem is mitigated. The residual noise leads to stochastic errors and also causes a small population control bias [48, 49, 22], which we control by increasing the walker number until undetectable levels of the bias are seen that are smaller than the stochastic error bars.
If both the box size $L$ and the momentum cutoff $k_{c}$ are large enough, the results obtained using the discretized model approach the continuum limit. We find that the energies converge exponentially in $L$, but only algebraically in $k_{c}$ (see Supplemental Material [50]).
Ground state energy – In the weak coupling limit, the ground state energy has been found analytically using perturbation theory [51, 33] and the path integral formalism [52]. For $P=0$, the ground state energy is
$\frac{E_{0}}{\hbar\omega_{\mathrm{LO}}}\approx-\alpha-0.06066\alpha^{2}-0.00844\alpha^{3}+\mathcal{O}(\alpha^{4}).$ (8)
While the numerical prefactors were rounded, this is exact to third order in $\alpha$. In the strong coupling limit, the ground state energy of Eq. (2) evaluates to
$\frac{E_{0}}{\hbar\omega_{\mathrm{LO}}}\approx-\frac{\alpha^{2}}{3}-0.955+\mathcal{O}(\alpha^{-2}).$ (9)
Figure 1: The ground state energy $E_{0}$ of the one-dimensional Fröhlich polaron, calculated using FCIQMC for various coupling strengths, compared with theoretical expansions in the weak [Eq. (8)] and strong coupling [Eq. (9)] regimes. The FCIQMC calculations used $L=6l_{0}$ and $k_{c}=12\pi/l_{0}$.
Figure 1 shows the transition from the weak coupling regime, through an intermediate coupling regime around $\alpha\approx 2$ where both expansions approximately overlap, into the strong coupling regime as the coupling strength $\alpha$ increases. While there is no phase transition separating the weak and strong coupling regimes, different physical pictures apply [53]. Our FCIQMC approach allows us to probe both regimes and the transition region comfortably. At $\alpha=2$ we estimate the discretization error to be around 1% and about an order of magnitude larger than the stochastic error [50].
Excited state energies – In order to calculate excited states numerically with FCIQMC, the computation is initialized with multiple starting vectors, which are then orthogonalized at every step (or a specified number of steps) of the computation using the Gram-Schmidt procedure. This ensures that the $n^{\text{th}}$ vector converges to the $n^{\text{th}}$ eigenvector, and the $n^{\text{th}}$ shift acts as an estimator for the $n^{\text{th}}$-lowest eigenvalue [19].
Even though the ground state calculation with the Hamiltonian (Bound excited states of Fröhlich polarons in one dimension) is free of the Monte Carlo sign problem, excited state calculations introduce the sign problem [28, 29] due to the necessary existence of nodes in the excited state wave functions [19]. The sign problem manifests itself at low walker number in fluctuating signs of all state vector coefficients and an average value of the shift $\langle S\rangle$ that is lower than the correct energy [17, 31]. Increasing the walker number above a critical value, the shift average stabilizes and becomes independent of the walker number, indicating that the sign problem is overcome. For the studied range of $\alpha$ values we find that $10^{7}$ walkers is sufficient, indicating that the sign problem is relatively weak in this case [31]. This is remarkable as the unconstrained phonon occupation makes the Hilbert space of the Fröhlich Hamiltonian (Bound excited states of Fröhlich polarons in one dimension) infinite. For details, see Appendix B in the End Matter.
Figure 2: The excitation energies of the first four excited states, with zero total momentum, calculated using FCIQMC for various coupling strengths and box sizes (error bars and symbols). The first bound excited state appears at $\alpha\approx 1.73$, in the intermediate coupling regime. The dashed horizontal line corresponds to the analytical result for the first discrete excitation energy at large $\alpha$. The shaded region shows the phonon continuum for $L\to\infty$. All states other than those labeled as odd are even under a parity transformation. Energies for $L=8l_{0}$ are only shown in the inset.
Figure 2 shows the excitation energies of a few low-lying excited states as function of $\alpha$ and for different box sizes $L$ at $P=0$. At the bottom of the phonon continuum (shaded region) there is always one state with an excitation energy of $1\hbar\omega_{\mathrm{LO}}$. It is given by $|\mathrm{C}\rangle=(\hat{a}_{0}^{\dagger}-\sqrt{{2\alpha l_{0}}/{L}})|\mathrm{GS}\rangle$, where $|\mathrm{GS}\rangle$ denotes the ground state. For weak coupling this is the first excited state in the spectrum. Due to the finite box size, the continuum of eigenstates with excitation energies above $1\hbar\omega_{\mathrm{LO}}$ is discretized. This discretizaton is most easily understood for vanishing coupling strength $\alpha=0$. In this case, the Hamiltonian is diagonal in the Fock state basis with eigenvalues ${E_{a,b}}/{\hbar\omega_{\mathrm{LO}}}=a+(2\pi bl_{0}/L)^{2}$, where $a\in\mathbb{N}$ is the total number of phonons, and $b\in\mathbb{Z}$ is the total phonon momentum in units of $\frac{2\pi\hbar}{L}$. As the coupling strength $\alpha$ increases, some of the excitation energies decrease.
As seen in Fig. 2, the first bound state with excitation energy below $1\hbar\omega_{\mathrm{LO}}$ appears at $\alpha\approx 1.73$, in the intermediate coupling regime. While the excitation energies in the discretized phonon continuum strongly depend on the box size $L$, this dependence disappears for the excited state below the continuum threshold. The value of $\alpha$ where the bound state emerges was found by repeating calculations with different $L$ at $\alpha=1.7$ and $\alpha=1.8$, which resulted in the consistent crossing point shown in the inset of Fig. 2. For $L=10l_{0}$, the threshold was found to be $\alpha=1.731\pm 0.003$.
The excitation energies were found to converge more slowly in $L$ than the ground state energy, showing that the excited states are larger in real space than the ground state. However, the excitation energies converge more quickly in the momentum cutoff $k_{c}$ than the absolute energies, and $k_{c}\approx 4\pi/l_{0}$ is large enough to find accurate excitation energies in the range of $\alpha$ reported here. See the Supplemental Material for more details [50].
Properties of bound excited states – The properties of the eigenstates can be studied more closely by looking at the density of virtual phonons in each mode, as shown in Fig. 3. The ground state gains more phonons as the coupling strength increases, with a larger density in the lower momentum modes. The continuum threshold state $|\mathrm{C}\rangle$ has increased phonon density only in the $k=0$ mode, as expected.
Figure 3: The density of virtual phonons in each mode $\langle j|a_{k}^{\dagger}a_{k}|j\rangle$ for the ground state $|j\rangle=|\mathrm{GS}\rangle$ and first two excited states, with $L=6l_{0}$ and $k_{c}=4\pi/l_{0}$, for various coupling strengths. The state $|\mathrm{C}\rangle$ is at the start of the continuum, and always has an excitation energy of $1\hbar\omega_{\mathrm{LO}}$. The state $|\mathrm{B}\rangle$ is a bound excited state for $\alpha\gtrsim 1.73$.
[…]
Spectral weight – The spectral weight of the state $|j\rangle$ at the total momentum $P$ from the Hamiltonian (Bound excited states of Fröhlich polarons in one dimension)
$Z_{j}^{(P)}=|\langle\mathrm{vac}|j\rangle|^{2},$ (10)
carries important information about the ability to detect the state in spectroscopic measurements (e.g. angle-resolved photoemission spectroscopy (ARPES) [54, 55, 56, 57] or polaron injection spectroscopy [58, 59, 60]). Figure 4 shows the spectral weight of the three lowest states in the discretized model. While the spectral weight of the ground state decreases $\sim 1-\alpha/2$ for small $\alpha$ as expected [8], the bound excited state carries a significant spectral weight $\approx 0.2$ near the threshold. The spectral weight of states within the continuum ($|\mathrm{C}\rangle$, and $|\mathrm{B}\rangle$ for $\alpha<1.73$) changes with box size and is expected to vanish in the limit $L\to\infty$. The spectral weight of the discrete states ($|\mathrm{GS}\rangle$, and $|\mathrm{B}\rangle$ for $\alpha>1.73$) is expected to depend weakly on the discretization of the model but a full convergence study is yet to be performed.
Figure 4: Spectral weight $Z_{j}^{(0)}$ of the three lowest energy states vs. the coupling strength $\alpha$ for $k_{c}=4\pi/l_{0}$ and two different box sizes. For $\alpha>1.73$ marked by the vertical dashed line the state $|\mathrm{B}\rangle$ becomes a bound excited state. The dotted line is the first order perturbation theory result $21-\alpha/$ 2 for the ground state.
Outlook – An outstanding question concerns the relevance of our findings to materials [5] or experimentally accessible synthetic quantum systems [61, 62]. The closest related candidate systems may be narrow quantum wires with strong surface optical phonons [63]. Recently predicted discrete excited phonon branches near a Mott-insulator–superfluid phase transition in a Bose-Hubbard model [64] may be a related phenomenon to the excited states studied in this work. An interesting avenue of future research will be to extend the FCIQMC approach to the three-dimensional Fröhlich model to find the threshold coupling constant for the bound excited state and independently check the diagrammatic Monte Carlo results of Ref. [8], which indicated the absence of bound excited states up to values of $\alpha\approx 8$. It will further be valuable to understand the class of problems in which the resource requirements for sign-problem amelioration by walker annihilation in projector quantum Monte Carlo decouple from the dimension of Hilbert space.
Acknowledgements.
Acknowledgements – We are grateful to Dmytro Kolisnyk for his help in working out the spectrum of the Hessian in the End Matter. This work was supported by the Marsden Fund of New Zealand (contract no. MAU2007) from government funding administered by the Royal Society Te Apārangi and by a summer scholarship from Te Whai Ao – Dodd-Walls Centre for Photonic and Quantum Technologies and the Physics Department, University of Auckland. We acknowledge support by the New Zealand eScience Infrastructure (NeSI) high-performance computing facilities in the form of a merit project allocation. Data availability – All numerical calculations were performed using the open source software Rimu.jl [27]. The data that support the findings of this article are openly available [65].